# The optimal hypercontractive constants for $\mathbb{Z}_{3}$

### v1

@@ 6.5. Estimates for $v_{t}$ and $\frac{v_{t}}{v}$
$v$ $=\frac{(1-t^{1-\alpha})(1-t^{1+\alpha})}{1+2t^{2}}=\frac{t}{1+2t^{2}}\left[t+\frac{1}{t}-\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)\right],$
$v_{t}$ $=\frac{-1}{1+2t^{2}}\left[\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)+\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)+\frac{2t}{1+2t^{2}}\right].$
@@
$\frac{\partial g}{\partial\alpha}$ $=\frac{-\log t}{1+2t^{2}}\left[\left(\frac{1}{\log t}+\frac{1-2t^{2}}{1+2t^{2}}\right)\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)+\alpha\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)\right]+2\alpha V^{\prime}_{1},$
$\frac{\partial^{2}g}{\partial\alpha^{2}}$ $=\frac{-(\log t)^{2}}{1+2t^{2}}\left[\left(\frac{2}{\log t}+\frac{1-2t^{2}}{1+2t^{2}}\right)\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)+\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)\right]+2V^{\prime}_{1},$
$\frac{\partial^{3}g}{\partial\alpha^{3}}$ $=\frac{-(\log t)^{3}}{1+2t^{2}}\left[\left(\frac{3}{\log t}+\frac{1-2t^{2}}{1+2t^{2}}\right)\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)+\alpha\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)\right]$
$=\frac{-(\log t)^{3}}{1+2t^{2}}\left(\frac{1}{t^{\alpha}}-t^{\alpha}\right)\left(-\frac{3}{\log t}-\frac{1-2t^{2}}{1+2t^{2}}+\alpha\frac{1+t^{2\alpha}}{1-t^{2\alpha}}\right).$
@@
$I:=A\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)-\left(t^{2\alpha}-\frac{1}{t^{2\alpha}}\right)+B\alpha\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)-4\alpha\log t,$
@@
$A=t+\frac{1}{t}-\left(t-\frac{1}{t}\right)\log t,\quad B=\left(t+\frac{1}{t}\right)\log t.$
@@
$\frac{\partial I}{\partial\alpha}$ $=(A\log t+B)\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)-\left(t^{2\alpha}+\frac{1}{t^{2\alpha}}\right)(2\log t)$
$\ \ \ \ +B\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)\log t-4\log t,$
$\frac{\partial^{2}I}{\partial\alpha^{2}}$ $=(A(\log t)^{2}+2B\log t)\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)-\left(t^{2\alpha}-\frac{1}{t^{2\alpha}}\right)(2\log t)^{2}$
$\ \ \ \ +B\alpha\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)(\log t)^{2},$
$\frac{\partial^{3}I}{\partial\alpha^{3}}$ $=(A(\log t)^{3}+3B(\log t)^{2})\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)-\left(t^{2\alpha}+\frac{1}{t^{2\alpha}}\right)(2\log t)^{3}$
$\ \ \ \ +B\alpha\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)(\log t)^{3},$
$\frac{\partial^{4}I}{\partial\alpha^{4}}$ $=(A(\log t)^{4}+4B(\log t)^{3})\left(t^{\alpha}-\frac{1}{t^{\alpha}}\right)-\left(t^{2\alpha}-\frac{1}{t^{2\alpha}}\right)(2\log t)^{4}$
$\ \ \ \ +B\alpha\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)(\log t)^{4}.$
@@
$=5\left(t+\frac{1}{t}\right)-\left(t-\frac{1}{t}\right)\log t-16\left(t^{\alpha}+\frac{1}{t^{\alpha}}\right)+\alpha\frac{1+t^{2\alpha}}{1-t^{2\alpha}}\left(t+\frac{1}{t}\right)(-\log t)$
$<5\left(t+\frac{1}{t}\right)-\left(t-\frac{1}{t}\right)\log t-32+\frac{1+t^{2}}{1-t^{2}}\left(t+\frac{1}{t}\right)(-\log t)$
$<5\times 5.2-\left(t-\frac{1}{t}\right)\log t-32+5.2\frac{1+t^{2}}{1-t^{2}}(-\log t)$
$=\frac{26}{5}\frac{1+t^{2}}{1-t^{2}}(-\log t)-\left(t-\frac{1}{t}\right)\log t-6<0.$
@@ Appendix B Univariate inequalities involving logarithms
$f_{2}(t)$ $:=\frac{(1+2t)^{2}}{4}f^{\prime}_{1}(t)$
$=\frac{(1-t)(24t^{6}-30t^{5}-78t^{4}-67t^{3}-75t^{2}-33t-65)}{36(1+2t^{2})^{2}}-\log t<0.$
@@ C.1. First proof
By symmetry, the curves $H(0,\cdot)$, $H(\alpha,\cdot)$, and $H(\alpha^{*},\cdot)$ intersect at a unique point. That is, there is a unique $t\in(0,1)$ such that
$H(0,t)=H(\alpha,\tfrac{1}{2})=H(-\alpha,\tfrac{1}{2}).$
By $\Phi(0,t)=(2,t)$ and (C.1), for the pairs $(p,2)$ and $(2,p^{*})$, the corresponding solutions of system (1.2) are respectively $(x_{1},y_{1})=(2^{\alpha-1},t)$ and $(x_{2},y_{2})=(t,2^{-\alpha-1}).$ By Theorem 1.1,

### v2

@@ 6.5. Estimates for $v_{t}$ and $\frac{v_{t}}{v}$
$v$ $=\frac{(1-t^{1-\alpha})(1-t^{1+\alpha})}{1+2t^{2}}=\frac{t}{1+2t^{2}}\Big[t+\tfrac{1}{t}-\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)\Big],$
$v_{t}$ $=\frac{-1}{1+2t^{2}}\left[\frac{1-2t^{2}}{1+2t^{2}}\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)+\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)+\frac{2t}{1+2t^{2}}\right].$
@@
$\frac{\partial g}{\partial\alpha}$ $=\frac{-\log t}{1+2t^{2}}\left[\left(\frac{1}{\log t}+\frac{1-2t^{2}}{1+2t^{2}}\right)\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)+\alpha\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)\right]+2\alpha V^{\prime}_{1},$
$\frac{\partial^{2}g}{\partial\alpha^{2}}$ $=\frac{-(\log t)^{2}}{1+2t^{2}}\left[\left(\frac{2}{\log t}+\frac{1-2t^{2}}{1+2t^{2}}\right)\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)+\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)\right]+2V^{\prime}_{1},$
$\frac{\partial^{3}g}{\partial\alpha^{3}}$ $=\frac{-(\log t)^{3}}{1+2t^{2}}\left[\left(\frac{3}{\log t}+\frac{1-2t^{2}}{1+2t^{2}}\right)\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)+\alpha\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)\right]$
$=\frac{-(\log t)^{3}}{1+2t^{2}}\left(\tfrac{1}{t^{\alpha}}-t^{\alpha}\right)\left(-\frac{3}{\log t}-\frac{1-2t^{2}}{1+2t^{2}}+\alpha\frac{1+t^{2\alpha}}{1-t^{2\alpha}}\right).$
@@
$I:=A\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)-\left(t^{2\alpha}-\tfrac{1}{t^{2\alpha}}\right)+B\alpha\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)-4\alpha\log t,$
@@
$A=t+\tfrac{1}{t}-\left(t-\tfrac{1}{t}\right)\log t,\quad B=\left(t+\tfrac{1}{t}\right)\log t.$
@@
$\frac{\partial I}{\partial\alpha}$ $=(A\log t+B)\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)-\left(t^{2\alpha}+\tfrac{1}{t^{2\alpha}}\right)(2\log t)+B\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)\log t-4\log t,$
$\frac{\partial^{2}I}{\partial\alpha^{2}}$ $=(A(\log t)^{2}+2B\log t)\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)-\left(t^{2\alpha}-\tfrac{1}{t^{2\alpha}}\right)(2\log t)^{2}+B\alpha\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)(\log t)^{2},$
$\frac{\partial^{3}I}{\partial\alpha^{3}}$ $=(A(\log t)^{3}+3B(\log t)^{2})\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)-\left(t^{2\alpha}+\tfrac{1}{t^{2\alpha}}\right)(2\log t)^{3}+B\alpha\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)(\log t)^{3},$
$\frac{\partial^{4}I}{\partial\alpha^{4}}$ $=(A(\log t)^{4}+4B(\log t)^{3})\left(t^{\alpha}-\tfrac{1}{t^{\alpha}}\right)-\left(t^{2\alpha}-\tfrac{1}{t^{2\alpha}}\right)(2\log t)^{4}+B\alpha\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)(\log t)^{4}.$
@@
$=5\left(t+\tfrac{1}{t}\right)-\left(t-\tfrac{1}{t}\right)\log t-16\left(t^{\alpha}+\tfrac{1}{t^{\alpha}}\right)+\alpha\frac{1+t^{2\alpha}}{1-t^{2\alpha}}\left(t+\tfrac{1}{t}\right)(-\log t)$
$<5\left(t+\tfrac{1}{t}\right)-\left(t-\tfrac{1}{t}\right)\log t-32+\frac{1+t^{2}}{1-t^{2}}\left(t+\tfrac{1}{t}\right)(-\log t)$
$<5\times 5.2-\left(t-\tfrac{1}{t}\right)\log t-32+5.2\frac{1+t^{2}}{1-t^{2}}(-\log t)$
$=\frac{26}{5}\frac{1+t^{2}}{1-t^{2}}(-\log t)-\left(t-\tfrac{1}{t}\right)\log t-6<0.$
@@
7. The biased two-point hypercontractivity: Proof of Theorem 1.6
Recall that the $\lambda$-biased Bernoulli random variable $X=X_{\lambda}$ $(0<\lambda<1)$ defined in (1.9):
$\mathbb{P}(X=1-\lambda)=\lambda,\quad\mathbb{P}(X=-\lambda)=1-\lambda.$
Then
$\|1+\rho X\|_{p}=\left(\lambda|1+(1-\lambda)\rho|^{p}+(1-\lambda)|1-\lambda\rho|^{p}\right)^{\frac{1}{p}}.$
We want to find the optimal constant $\sigma_{p,q}(\lambda)\in[0,1]$ such that
$\|1+r\rho X\|_{q}\leq\|1+\rho X\|_{p}\quad\text{for all $0\leq r\leq\sigma_{p,q}(\lambda)$ and all $\rho\in\mathbb{R}$}.$
We only need to consider $\rho$’s such that $1+\rho X\geq 0$. Thus we only need to consider
$\Big(\lambda(1+(1-\lambda)r\rho)^{q}+(1-\lambda)(1-\lambda r\rho)^{q}\Big)^{\frac{1}{q}}\leq\Big(\lambda(1+(1-\lambda)\rho)^{p}+(1-\lambda)(1-\lambda\rho)^{p}\Big)^{\frac{1}{p}}$
with
(7.1) $-\frac{1}{1-\lambda}\leq\rho\leq\frac{1}{\lambda}.$
As in (1.15), we define the defect function
$G(\lambda,r,\rho):=G(\lambda,p,q,r,\rho)=$ $\Big(\lambda(1+(1-\lambda)\rho)^{p}+(1-\lambda)(1-\lambda\rho)^{p}\Big)^{\frac{1}{p}}$
- $\Big(\lambda(1+(1-\lambda)r\rho)^{q}+(1-\lambda)(1-\lambda r\rho)^{q}\Big)^{\frac{1}{q}}$
with $r\in(0,1)$ and $\rho$ satisfying (7.1).
Lemma 7.1.
For $\lambda\in(0,1/2)$ and $1<p<q<\infty$, one has $0<\sigma_{p,q}(\lambda)<\sqrt{(p-1)/(q-1)}.$
Proof.
The proof is almost the same as that for Lemma 3.1. We only need to take the Taylor expansion of $G(\lambda,r,\rho)$ at $\rho=0$:
$G=\tfrac{\lambda(1-\lambda)}{2}\Big[(p-1)-(q-1)r^{2}\Big]\rho^{2}+\tfrac{\lambda(1-\lambda)(1-2\lambda)}{6}\Big[(p-1)(p-2)-(q-1)(q-2)r^{3}\Big]\rho^{3}+O(\rho^{4}).$
The remaining proof is the same as that in Lemma 3.1. ∎
Lemma 7.2.
For $\lambda\in(0,1)\setminus\{1/2\}$ and $1<p<q<\infty$, one has
$G(\lambda,\sigma_{p,q}(\lambda),\tfrac{1}{\lambda})>0\quad\text{and}\quad G(\lambda,\sigma_{p,q}(\lambda),-\tfrac{1}{1-\lambda})>0.$
Proof.
By the definition of $\sigma:=\sigma_{p,q}(\lambda)$, we have $G(\lambda,\sigma,\frac{1}{\lambda})\geq 0$. Now we prove that
$G(\lambda,\sigma,\tfrac{1}{\lambda})>0.$
Assume by contradiction that $G(\lambda,\sigma,\tfrac{1}{\lambda})=0$. Then the left derivative of $G(\lambda,\sigma,\cdot)$ at $\rho=\tfrac{1}{\lambda}$ is non-positive. That is,
$G(\lambda,\sigma,\tfrac{1}{\lambda})=0\quad\text{and}\quad\tfrac{\partial G}{\partial\rho}(\lambda,\sigma,(\tfrac{1}{\lambda})^{-})\leq 0.$
Directive computation gives
$\left\{\begin{aligned} &\lambda^{\frac{q}{p}-q}=\lambda\left(1+\tfrac{1-\lambda}{\lambda}\sigma\right)^{q}+(1-\lambda)(1-\sigma)^{q},\\ &\lambda^{\frac{q}{p}-q}\leq\sigma\left[\left(1+\tfrac{1-\lambda}{\lambda}\sigma\right)^{q-1}-(1-\sigma)^{q-1}\right].\end{aligned}\right$.
It follows that
$\lambda\left(1+\tfrac{1-\lambda}{\lambda}\sigma\right)^{q}+(1-\lambda)(1-\sigma)^{q}\leq\sigma\left[\left(1+\tfrac{1-\lambda}{\lambda}\sigma\right)^{q-1}-(1-\sigma)^{q-1}\right],$
and hence
$\left(1+\tfrac{1-\lambda}{\lambda}\sigma\right)^{q-1}\lambda(1-\sigma)+(1-\sigma)^{q-1}[(1-\lambda)(1-\sigma)+\sigma]\leq 0.$
This is clearly impossible, since $\lambda,\sigma\in(0,1)$. Therefore $G(\lambda,\sigma_{p,q}(\lambda),\tfrac{1}{\lambda})>0$.
Notice that
$G(\lambda,r,\rho)=G(1-\lambda,r,-\rho)\quad\text{and}\quad\sigma_{p,q}(\lambda)=\sigma_{p,q}(1-\lambda).$
Combing the previous discussion, one has
$G(\lambda,\sigma_{p,q}(\lambda),-\tfrac{1}{1-\lambda})=G\left(1-\lambda,\sigma_{p,q}(1-\lambda),\tfrac{1}{1-\lambda}\right)>0.$
This completes the whole proof. ∎
Lemma 7.3.
There exists $\rho_{0}\in(-\frac{1}{1-\lambda},\frac{1}{\lambda})\setminus\{0\}$ such that
$G(\lambda,\sigma_{p,q}(\lambda),\rho_{0})=0\quad\text{and}\quad\tfrac{\partial G}{\partial\rho}(\lambda,\sigma_{p,q}(\lambda),\rho_{0})=0.$
Proof.
Using Lemmas 7.1 and 7.2, the proof is exactly the same as Proposition 4.1. ∎
Proof of Theorem 1.6.
By Lemma 7.3, the pair $(r,\rho)=(\sigma_{p,q}(\lambda),\rho_{0})$ satisfies the following system:
$\left\{\begin{aligned} &\Big(\lambda(1+(1-\lambda)\rho)^{p}+(1-\lambda)(1-\lambda\rho)^{p}\Big)^{\frac{1}{p}}=\Big(\lambda(1+(1-\lambda)r\rho)^{q}+(1-\lambda)(1-\lambda r\rho)^{q}\Big)^{\frac{1}{q}},\\ &\rho\frac{(1+(1-\lambda)\rho)^{p-1}-(1-\lambda\rho)^{p-1}}{\lambda(1+(1-\lambda)\rho)^{p}+(1-\lambda)(1-\lambda\rho)^{p}}=r\rho\frac{(1+(1-\lambda)r\rho)^{q-1}-(1-\lambda r\rho)^{q-1}}{\lambda(1+(1-\lambda)r\rho)^{q}+(1-\lambda)(1-\lambda r\rho)^{q}}.\end{aligned}\right$.
Define the following change of variables:
(7.2) $x=\frac{1-\lambda\rho}{1+(1-\lambda)\rho},\quad y=\frac{1-\lambda r\rho}{1+(1-\lambda)r\rho}.$
When
$0<r<1,\quad-\frac{1}{1-\lambda}<\rho<\frac{1}{\lambda},$
the range of $(x,y)$ is
[…]
Then $(x,y)$ satisfies the following equation system:
$\left\{\begin{aligned} &\frac{\big(\lambda+(1-\lambda)x^{p}\big)^{\frac{1}{p}}}{\lambda+(1-\lambda)x}=\frac{\big(\lambda+(1-\lambda)y^{q}\big)^{\frac{1}{q}}}{\lambda+(1-\lambda)y},\\ &\frac{(1-x)(1-x^{p-1})}{\lambda+(1-\lambda)x^{p}}=\frac{(1-y)(1-y^{q-1})}{\lambda+(1-\lambda)y^{q}},\end{aligned}\right$.
which is the desired equation-system (1.10). Reversing the change-of-variable (7.2), one obtains
$\sigma_{p,q}(\lambda)=\frac{1-y}{1-x}\cdot\frac{\lambda+(1-\lambda)x}{\lambda+(1-\lambda)y},$
as desired. ∎
@@ Appendix B Univariate inequalities involving logarithms
$f_{2}(t)$ $:=\frac{(1+2t)^{2}}{4}f^{\prime}_{1}(t)=\frac{(1-t)(24t^{6}-30t^{5}-78t^{4}-67t^{3}-75t^{2}-33t-65)}{36(1+2t^{2})^{2}}-\log t<0.$
@@ C.1. First proof
By symmetry, the curves $H(0,\cdot)$, $H(\alpha,\cdot)$, and $H(\alpha^{*},\cdot)$ intersect at a unique point. That is, there is a unique $t\in(0,1)$ such that $H(0,t)=H(\alpha,\tfrac{1}{2})=H(-\alpha,\tfrac{1}{2}).$ By $\Phi(0,t)=(2,t)$ and (C.1), for the pairs $(p,2)$ and $(2,p^{*})$, the corresponding solutions of system (1.4) are respectively $(x_{1},y_{1})=(2^{\alpha-1},t)$ and $(x_{2},y_{2})=(t,2^{-\alpha-1}).$ By Theorem 1.2,